OpenStudy (anonymous):

Find a quadratic equation with roots -1+4i and -1-4i

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you wanna do it the real quick way?

OpenStudy (misty1212):

or the slow way?

OpenStudy (anonymous):

Real quick please!

OpenStudy (misty1212):

ok it is a matter of knowing that is \(a+bi\)is the root of a quadratic, it is \[x^2-2ax+(a^2+b^2)\]

OpenStudy (misty1212):

in your case \(a=-1,b=4\) so you get \[x^2-2\times (-1)x+(-1)^2+4^2=x^2+2x+17\]

OpenStudy (misty1212):

pretty quick huh? you do have to memorize that though, there is a slower way to do it

OpenStudy (anonymous):

So you just plug in the numbers into the equation? That's not bad at all

OpenStudy (misty1212):

no it is not bad make sure you know what \(a\) and \(b\) are , there is no \(i\) involved just \[\huge x^2-2ax+a^2+b^2\]

OpenStudy (misty1212):

bet $5 your teacher didn't mention it did they?

OpenStudy (anonymous):

Haha no not at all! Thank you very much. I appreciate your help

OpenStudy (misty1212):

\[\color\magneta\heartsuit\]